Night Report 2016-07-30

Observing: Morgan Fraser, Tom Reynolds

Support Team: Seppo Mattila

TAT: Lluís Galbany, Stefan Taubenberger

Weather conditions

End of night: Clouds caused closure, but reopened. Seeing ~1"

------------------------------------------------

Start of night: Some thin cloud, seeing ~1"

Middle of night: Thin clouds at horizon, seeing ~1"

23:33 End of twilight

00:06 AT2016eiy (follow-up): Gr11+Gr16, slit 1.0", 900s each. Mis-acquired initially, took spectrum of nearby bright source, lost ~40 mins

01:35 SN2016adj (follow-up): Attempted to acquire for spectrum, could not reliably find object. Took 3x100s i-band imaging to determine if object still present

01:57 Standard, LTT7379

02:16 LSQ15adm (follow-up): Bgi 9x200s, Vr 6x200s. Small software fault cost 5-10 minutes

05:06 Attempted image analysis as there was evidence of PSF distortion in final LSQ15adm images. Actuator problem caused downtime for ~40 minutes

05:48 OGLE15sd (follow-up): 12x200s r

06:40 ASASSN16hw (classification): SN Ia, pre max, approx -5 days

06:49 AT2016egz (follow-up): Gr13, slit 1.0", 1200s (also refocused)

07:20 OGLE15qz (follow-up): 9x200s g

08:00 PSNJ2150 (follow-up): took spectrum of what we thought was target, was nearby source. Target was very faint, not detected in acquisition image

09:15 CLOSED DUE TO CLOUDS

09:55 REOPENED

10:00 SN2009ip (follow-up): 2x100s r, 10x150s i

10:06 Start of twilight

10:45 Standard, VMA2, only 1" slit

Quite cloudy, unable to take skyflats

Calibrations

Evening: EFOSC biases, spec flats

------------------------------------------------

Start of night: Had to restart TCS, prevented us from taking skyflats

22:55 Image analysis

23:13 SN2016ase (follow-up): Gr13, slit 1.0", 2700